OpenStudy (anonymous):

simplify; 2sqrt{2}-3sqrt{3}+sqrt{27}-sqrt{50} A)sqrt{-3} B)-3sqrt{2} C)7sqrt{2}-6sqrt{3} D)2sqrt{2}-2sqrt{5}

OpenStudy (anonymous):

\[2\sqrt{2}-3\sqrt{3}+\sqrt{27}-\sqrt{50}\] \[=2\sqrt{2}-3\sqrt{3}+3\sqrt3-5\sqrt2\] \[=-3\sqrt{2}\] ans=B

OpenStudy (anonymous):

B

OpenStudy (sriram):

2{2}^1/2 -3{3}^1/2 +3{3}^1/2-5{2}^1/2 2{2}^1/2 -5{2}1/2 {2}^1/2 *[2-5] =-3*{2}1/2